OpenStudy (anonymous):

Factor completely: 2x3 + 10x2 + 4x + 20

mathslover (mathslover):

first of all find the factors of +20 do you know the factors of 20?

OpenStudy (callisto):

Take out the common factor of the first two terms and last two terms. 2x^3 + 10x2 + 4x + 20 = 2x^2 (x+5) + 4(x+5) = ...?

OpenStudy (anonymous):

2x^3 + 10x^2 + 4x + 20 = 0 (2x+10)(x^2+2) = 0 2x = -10 ; x^2 = -2 x = -5 ; x = √ -2 x = -5 ; x = + or -√ 2

OpenStudy (callisto):

@ashna It's factoring the expression, not solving equation. Moreover, please don't give direct answers to the question. You may refer to our Code of Conduct at httlp://www.opnstudy.com/code-of-conduct for more information. Thanks.

OpenStudy (anonymous):

okay .. sorry !

mathslover (mathslover):

@Callisto it can be done in this way also : factors of 20 = \(\large{\pm 1,\pm 2,\pm 4,\pm 5,\pm 10, \pm20}\) now given p(x) = 2x3 + 10x2 + 4x + 20 p(-5)=-250+250-20+20=0 hence -5 is a zero of the polynomial p(x) that is (x+5) is a factor of p(x)

mathslover (mathslover):

M i right in this way also ? please dont consider this as direct answer... it is my confusion it may be wrong and also the answer is not yet declared

OpenStudy (callisto):

I haven't learnt this way. It may work, but for this question, it's not necessary to use this way, I think. @mathslover

mathslover (mathslover):

ok wait lemme show my full work

OpenStudy (callisto):

Well, I understand this method. Probably, you're doing by factor theorem. It works but it just takes a long time for this question as you have to test for many factors.

mathslover (mathslover):

this is what i did

mathslover (mathslover):

Very right @Callisto It is factor theorem .. but yes it takes a long time

mathslover (mathslover):

thanks to god that it was 20 .. if it was some other large number like 4000993 then it was next to impossible to do by this

OpenStudy (callisto):

For this type of problem,usually, I would see if it can be done by taking out factors/identities first. If not, then I have to use factor theorem.

mathslover (mathslover):

right